OpenStudy (anonymous):

Solve 5 – 3x = 8 -13/3 13/3 –1 1

OpenStudy (anonymous):

isolate x dude

OpenStudy (anonymous):

try rewriting the equation as -3x+5=8

OpenStudy (lilg132):

or try bringing the 5 to the other side and dividing everything by 3

OpenStudy (anonymous):

how u do that? I just came from cuba and im trying to pass the 1th semester and

OpenStudy (anonymous):

Ok we'll do this step by step. First, copy the original equation: 5-3x=8 Next, get x on it's own. So we need to get rid of the 5 on the left side. But whatever you do on one side you need to do to the other. Subtracting 5 on both sides gives us -3x=3 Finally, divide by -3. Remember, the -3x is really -3 times x, so we divide to remove it. So, the answer is x equals 3 divided by -3.

OpenStudy (anonymous):

Does that answer your question?

OpenStudy (lilg132):

5 - 3x = 8 if you bring 5 to the other side it wont be a positive number anymore it will change to a negative so -3x = 8 - 5 which is -3x = 3 now divide both sides by 3 because you want to get x on its own -3x/3 = 3/3 which gives you -x = 1 now swap them around to get x instead of -x and the answer is -1 = x